OpenStudy (anonymous):

multiply: (8a-8)(9a^2+7a+5)

OpenStudy (mathmagician):

(8a-8)(9a^2+7a+5)=72a^3+56a^2+40a-72a^2-56a-40=72a^3-16a^2-16a-40

OpenStudy (anonymous):

i got 72a^3-72a^2+40a-40

OpenStudy (anonymous):

\[ \begin{align} &&9a^2 + 7a &+ 5 \\ \times &&8a &- 8\\ \hline\\ &&-72a^2-56a &-40\\ &&72a^3+56a^2+40a\\ \hline\\ &&72a^3-16a^2-16a &-40\\ \end{align}\]

OpenStudy (anonymous):

thanks